alones (alones):

What is the distance between the points (2, 5) and (5, 7)?

alones (alones):

@ShadowLegendX heelp mee

OpenStudy (erak):

Does the distance formula sound familiar? d = sqrt((5-2)^2 + (7-5)^2))

alones (alones):

You mean this d =\(\sqrt(5-2)^2 + (7-5)^2)\)

OpenStudy (erak):

yeah

RhondaSommer (rhondasommer):

so solve dat ^

RhondaSommer (rhondasommer):

5-2=?

alones (alones):

3

OpenStudy (erak):

basically you'll have d = sqrt(3^2 + 2^2) d = sqrt(9+4) d = sqrt(13) units

alones (alones):

yeah i got \[\sqrt{13}\]

RhondaSommer (rhondasommer):

correct